How much of the Aztec coffee, which sells for $6.00 per pound, and how much of he Cinnamon coffee, which sells for $8.00 per pound, should be mixed to make a 5 lb of Winterfresh blend to be sold at $7.20 per pound?
Answer by mananth(16946)

Price ---------------- quantity
Aztec coffee $6.00 ---------------- x pounds
Cinnamon coffee $8.00 ------ 5 - x pounds
Mixture $7.20 ---------------- 5
5
6 x + 8 ( 5 - x ) = 36
6 x + 40 - 8 x = 36
6 x - 8 x = 36 - -40
-2 x = -4
/ -2
x = 2 pounds $6.00 Aztec coffee
3 pounds $8.00 Cinnamon coffee

m.ananth@hotmail.ca